1A. The manager of the local emu farm reports that only 34% of the emus are hostile to humans. If the farm has 250 emus, how many emus are hostile to humans?

B. The manager of the local emu farm reports that only 34% of the emus are hostile to humans. If the farm has 250 emus, how many emus are not hostile to humans?

C. The manager of the emu farm in Madera county reports the same proportion of 34% of hostile emus. If there are 132 emus that are nice to humans in the Madera county farm, what is the total emu population of that farm?
Answer #1

1A) Total = 250

p(hostile) = 0.34

Total hostile = 250* 0.34 = 85

1B) Total not hostile = 250 - 85 = 165

1C) total nice = 132

P(nice) = 1 - 0.34 = 0.66

Total nice = Total * P(nice)

Total = (total nice) / P(nice) = 132/0.66 = 200

Hence there are total 200 emu population of that farm.

Please comment if any doubt. Thank you.